# Smooth polynomials with several prescribed coefficients

### v2

@@ 7. Proof of Theorem 3
(33) $\begin{split}M_{2}(r,k)&=\left(q^{n-\ell-r-k}\rho\left(u\right)\left(u\log u\right)^{1/8}\kappa\left(1+\log_{q}\kappa\right)+R(r,k)\right)\\ &\qquad\cdot\sum_{\begin{subarray}{c}\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}\end{subarray}}\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{g_{0}t^{k}}\right).\end{split}$
@@
Then we have by Lemmas 9 and 26 that
$\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{g_{0}t^{k}}\right)$
$=$ $\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}}\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\end{subarray}}\mathbf{e}\left(-\frac{af}{g_{0}t^{k}}\right)$
$=$ $\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ \gcd(g_{0},t)=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{1}|<|g_{0}|\\ \gcd(a_{1},g_{0})=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{2}|<|t^{k}|\\ \gcd(a_{t},t)=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\end{subarray}}\mathbf{e}\left(-\frac{a_{1}f}{g_{0}}-\frac{a_{2}f}{t^{k}}\right)$
$=$ $\sum_{\begin{subarray}{c}|c|<|t^{k}|\\ \gcd(c,t)=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{2}|<|t^{k}|\\ \gcd(a_{t},t)=1\end{subarray}}\mathbf{e}\left(-\frac{a_{2}c}{t^{k}}\right)\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ \gcd(g_{0},t)=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{1}|<|g_{0}|\\ \gcd(a_{1},g_{0})=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv c\bmod t^{k}\end{subarray}}\mathbf{e}\left(-\frac{a_{1}f}{g_{0}}\right)$
$=$ $\sum_{\begin{subarray}{c}|c|<|t^{k}|\\ \gcd(c,t)=1\end{subarray}}\mu(t^{k})\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ \gcd(g_{0},t)=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{1}|<|g_{0}|\\ \gcd(a_{1},g_{0})=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv c\bmod t^{k}\end{subarray}}\mathbf{e}\left(-\frac{a_{1}f}{g_{0}}\right)$
$\ll$ $|\mu(t^{k})|\frac{q^{n}}{q^{\#{\mathcal{I}}}}q^{2rC_{1}(\#{\mathcal{I}}+k)/n}.$
where $C_{1}$ as in Lemma 26. Thus by (33)
(34) $\begin{split}M_{2}(r,k)&\ll\left(q^{n-\ell-r-k}\rho\left(u\right)\left(u\log u\right)^{1/8}(r+k)\left(1+\log_{q}(r+k)\right)+R(r,k)\right)\\ &\quad\cdot|\mu(t^{k})|\frac{q^{n}}{q^{\#{\mathcal{I}}}}q^{2rC_{1}(\#{\mathcal{I}}+k)/n}.\end{split}$
@@
$C_{2}=1-4C_{1}\frac{\#{\mathcal{I}}+1}{n}.$
As #$\#{\mathcal{I}}<n/10$, we have then $C_{2}\geq 1/5$ and $22C_{1}(\#I+1)/n\leq 1/$ 2. As $M_{2}(r,k)=0$ for $k\geq 2$, we have by (32) and (34) that
(35) $\begin{split}M_{2}&\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(u\right)\left(u\log u\right)^{1/8}\sum_{r\geq K}\frac{r\left(1+\log_{q}r\right)}{q^{C_{2}r}}+O\left(q^{-\frac{(1-4\varepsilon)n+m}{4}+2}e^{\varepsilon(n-m)/4}\right)\right)\\ &\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(u\right)\left(u\log u\right)^{1/8}\frac{K\left(1+\log_{q}K\right)}{q^{C_{2}K}}+O\left(q^{-\frac{(1-4\varepsilon)n+m}{4}+2}e^{\varepsilon(n-m)/4}\right)\right)\\ &\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(u\right)\left(u\log u\right)^{1/8}\frac{K\left(1+\log_{q}K\right)}{q^{C_{2}K}}+O\left(q^{-\frac{(1-4\varepsilon)n+m}{4}+2}e^{\varepsilon(n-m)/4}\right)\right).\end{split}$
@@
(36) $\begin{split}\int_{{\mathfrak{M}}}S(\xi;n,m)\overline{S_{\mathcal{J}}(\xi)}\mathrm{d}\xi&=\frac{\Psi(n,m)}{q^{\#I}}\left(1+O\left(\left(u\log u\right)^{1/8}\frac{\delta^{-1}\log\delta}{q^{C/\delta}}+\frac{1}{q^{i_{0}+1}}\frac{u^{1/2}(\log u)^{5/2}}{m}\right)\right)\\ &\quad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda_{0}(n,m,I,u)+\Lambda_{1}(n,m,I,u)+O\left(q^{-\frac{(1-8\varepsilon)n+2m}{8}+2}e^{\varepsilon\frac{n-m}{2}}\right)\right).\end{split}$
@@
By (27), we get in both cases
$S(\xi,m,n)\ll m^{2}n^{1/2}q^{(7n+m+4)/8}.$
@@
(37) $\begin{split}\int_{{\mathfrak{m}}}S(\xi;n,m)\overline{S_{\mathcal{J}}(\xi)}\mathrm{d}\xi&\ll m^{2}n^{1/2}q^{(7n+m+4)/8}\int_{{\mathfrak{m}}}|S_{\mathcal{J}}(\xi)|\mathrm{d}\xi\\ &\ll m^{2}n^{1/2}q^{(7n+m+4)/8}\int_{\mathbf{T}}|S_{\mathcal{J}}(\xi)|\mathrm{d}\xi\\ &\ll m^{2}n^{1/2}q^{(7n+m+4)/8}.\end{split}$
By replacing $\varepsilon$ if needed in (36), Theorem 3 follows from (36)and (37).
@@ 8. Proof of Theorem 1
Theorem 1 follows directly form Theorem 3. Indeed, if $0\in{\mathcal{I}}$ and $\alpha_{0}\neq 0$, then $\Lambda_{0}(n,m,I,u)=0$ and
$\Lambda_{1}(n,m,I,u)=-\frac{1}{m}\frac{1}{q-1}\rho^{\prime}\left(u\right)\sim\frac{1}{m}\frac{1}{q}\rho\left(u\right)\log u$
by Lemmas 12 and 13, which is absorbed by the error term.
9. Proof of Corollaries 2 and 4
@@
For Corollary 4, if $q\rightarrow\infty$, then
$|\Lambda_{0}(n,m,I,u)|,|\Lambda_{1}(n,m,I,u)|\leq c(n,m)\frac{1}{q}$
for some constant $c(n,m)$ depends only on $n$ and $m$ (and thus on $u$).
If $n\rightarrow\infty$ , then by Lemmas 12 and 13,
$|\rho^{\prime}((n-i_{j})/m)|\leq c\rho(u)u^{1/8}(\log u)^{9/8}$
for some absolute constant $c$, and as $\nu\leq\#{\mathcal{I}}=n\delta$, thus
$|\Lambda_{0}(n,m,I,u)|\leq c(q)\rho(u)\frac{u^{1/8}(\log u)^{9/8}}{m}n\delta=c(q)\rho(u)u^{9/8}(\log u)^{9/8}\delta=o(\rho(u)).$
Similarly,
$|\Lambda_{1}(n,m,I,u)|\leq c_{1}(q)\frac{1}{m}\rho^{\prime}\left(\frac{n-i_{\nu}}{m}\right)\leq c_{2}(q)\rho(u)\frac{u^{1/8}(\log u)^{9/8}}{m}=o(\rho(u)).$
@@ Acknowledgment
The author wishes to thank Ofir Gorodetsky and Eugenijus Manstavic̆ius for useful discussions and valuable comments. He also thanks the anonymous referee to bring the attention to [1] which allowed to improve Lemma 16.
The author was was partially supported by NRDI (National Research Development and Innovation Office, Hungary) grant FK 142960,

### v3

@@ 7. Proof of Theorem 4
(44) $\begin{split}M_{2}(r,k)&=\left(q^{n-\ell-(r+k)/5}\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)e^{T(n,m)}(r+k)\left(1+\log_{q}(r+k)\right)+R(r,k)\right)\\ &\qquad\cdot\sum_{\begin{subarray}{c}\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}\end{subarray}}\left|\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}S_{\mathcal{J}}\left(\frac{a}{g_{0}t^{k}}\right)\right|.\end{split}$
@@
Then we have by Lemmas 11 and 27 that
$\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}}\left|\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}S_{\mathcal{J}}\left(\frac{a}{g_{0}t^{k}}\right)\right|$
$=$ $\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}}\left|\sum_{\begin{subarray}{c}|a_{1}|<|g_{0}|\\ \gcd(a_{1},g_{0})=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{2}|<|t^{k}|\\ \gcd(a_{2},t)=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\end{subarray}}\mathbf{e}\left(-\frac{a_{1}f}{g_{0}}-\frac{a_{2}f}{t^{k}}\right)\right|$
$=$ $\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}}\left|\sum_{\begin{subarray}{c}|c|<|t^{k}|\\ \gcd(c,t)=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{2}|<|t^{k}|\\ \gcd(a_{2},t)=1\end{subarray}}\mathbf{e}\left(-\frac{a_{2}c}{t^{k}}\right)\sum_{\begin{subarray}{c}|a_{1}|<|g_{0}|\\ \gcd(a_{1},g_{0})=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv c\bmod t^{k}\end{subarray}}\mathbf{e}\left(-\frac{a_{1}f}{g_{0}}\right)\right|$
$=$ $\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}}\left|\sum_{\begin{subarray}{c}|c|<|t^{k}|\\ \gcd(c,t)=1\end{subarray}}\mu(t^{k})\sum_{\begin{subarray}{c}|a_{1}|<|g_{0}|\\ \gcd(a_{1},g_{0})=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv c\bmod t^{k}\end{subarray}}\mathbf{e}\left(-\frac{a_{1}f}{g_{0}}\right)\right|$
(45) $=$ $|\mu(t^{k})|\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}}\left|\sum_{\begin{subarray}{c}|c|<|t|\\ \gcd(c,t)=1\end{subarray}}\mu(t^{k})\sum_{\begin{subarray}{c}|a_{1}|<|g_{0}|\\ \gcd(a_{1},g_{0})=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv c\bmod t^{k}\end{subarray}}\mathbf{e}\left(-\frac{a_{1}f}{g_{0}}\right)\right|.$
Clear, (7) is zero, unless,
(46) $k\leq 1.$
If $k=0$, we have by Lemma 27 that (7) is
(47) $O\left(\frac{q^{n}}{q^{\#{\mathcal{I}}}}q^{2rC_{1}\#{\mathcal{I}}/n}\right).$
If $k=1$, for all $|c|\leq|t|$, we can add the constrains to $\mathcal{J}$} that the constant term of $f$ is $c$ to derive that (7) is
(48) $O\left(\frac{q^{n}}{q^{\#{\mathcal{I}}}}q^{2rC_{1}(\#{\mathcal{I}}+1)/n}\right).$
Thus (47) and (48) yield that by (44) we have
(49) $\begin{split}M_{2}(r,k)&\ll\Bigg{(}q^{n-\ell-(r+k)/5}\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)e^{T(n,m)}(r+k)\left(1+\log_{q}(r+k)\right)+R(r,k)\Bigg{)}\\ &\quad\cdot|\mu(t^{k})|\frac{q^{n}}{q^{\#{\mathcal{I}}}}q^{2rC_{1}(\#{\mathcal{I}}+k)/n}.\end{split}$
@@
$C_{2}=\frac{1}{5}-2C_{1}\frac{\#{\mathcal{I}}+1}{n}.$
As ${\mathcal{I}}<n/24$, we have $2C_{1}(\#{\mathcal{I}}+1)/n\leq 1/6$ and thus $C_{2}\geq 1/30$. Then we have by (42), (43),(46) and (49) that
(50) $\begin{split}M_{2}&\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)e^{T(n,m)}\sum_{r\geq K}\frac{r\left(1+\log_{q}r\right)}{q^{C_{2}r}}+O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\frac{1}{2}+\ell}e^{\varepsilon(\ell+\kappa)}\right)\right)\\ &\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)e^{T(n,m)}\frac{K\left(1+\log_{q}K\right)}{q^{C_{2}K}}+O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\frac{1}{2}+\ell}e^{\varepsilon(\ell+\kappa)}\right)\right).\end{split}$
@@
(51) $\begin{split}\int_{{\mathfrak{M}}}S(\xi;n,m)\overline{S_{\mathcal{J}}(\xi)}\mathrm{d}\xi&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+O\left(e^{T(n,m)}\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\right)\right)\\ &\quad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}})+O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\ell+\nu+\frac{1}{2}}e^{\varepsilon(\ell+\kappa)}\right)\right).\end{split}$
As $\nu\leq i_{\nu}<\kappa$, we get by choosing
(52) $\ell=\kappa+1,$
that the error term here is
(53) $O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\ell+\nu+\frac{1}{2}}e^{\varepsilon(\ell+\kappa)}\right)=O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+3\kappa+\frac{3}{2}}e^{2\varepsilon\kappa}.\right)$
@@
We get in both cases by (14) and (52) that
$S(\xi,m,n)\ll m^{2}n^{1/2}q^{n-\kappa/2}.$
@@
(54) $\begin{split}\int_{{\mathfrak{m}}}S(\xi;n,m)\overline{S_{\mathcal{J}}(\xi)}\mathrm{d}\xi&\ll m^{2}n^{1/2}q^{n-\kappa/2}\int_{{\mathfrak{m}}}|S_{\mathcal{J}}(\xi)|\mathrm{d}\xi\\ &\ll m^{2}n^{1/2}q^{n-\kappa/2}\int_{\mathbf{T}}|S_{\mathcal{J}}(\xi)|\mathrm{d}\xi\\ &\ll m^{2}n^{1/2}q^{n-\kappa/2}.\end{split}$
By replacing $\varepsilon$ if needed in (51), Theorem 4 follows from (51), (53) and (54).
@@ 8. Proof of Theorem 1
Theorem 1 follows immediately from Theorem 4 by noting that $\Lambda(n,m,{\mathcal{I}})=0$ if $0\in{\mathcal{I}}$ and $\alpha_{0}\neq 0$.
9. Proof of Corollaries 3 and 6
@@
For Corollary 6, consider $\Lambda(n,m,{\mathcal{I}},\kappa)$. For its terms we have
$\frac{1}{q^{n-j-1}}\left(\Psi(n-i_{j}-1,m)-\frac{1}{q}\Psi(n-i_{j},m)\right)\ll\frac{1}{q^{i_{j}-j}}\left(\rho\left(\frac{n-i_{j}-1}{m}-\frac{n-i_{j}-1}{2m^{2}}\right)\right).$
As $0\not\in{\mathcal{I}}$, we have $i_{j}\geq j+1$, thus
$\Lambda(n,m,{\mathcal{I}},\kappa)\ll\frac{1}{q}$
provided, that $m$ and $n$ are fixed.
@@ Acknowledgment
The author wishes to thank Ofir Gorodetsky and Eugenijus Manstavic̆ius for useful discussions and valuable comments. He also thanks the anonymous referee for calling attention to [1], which allowed us to improve Lemma 17.
The author was partially supported by NRDI (National Research Development and Innovation Office, Hungary) grant FK 142960 and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.